Determine the length of arc of curve 𝑦

2 = 2π‘₯ (π‘“π‘œπ‘Ÿ 0 ≀ π‘₯ ≀

1

2

) in the firstΒ 

quadrant.


1
Expert's answer
2021-05-07T11:47:51-0400

Determine the length of arc of curve "\ud835\udc66^2 = 2\ud835\udc65" (π‘“π‘œπ‘Ÿ 0 ≀ π‘₯ ≀ 12)

"y=f(x)=\sqrt{2x} \newline
y'=\sqrt{2}\frac{1}{2}x^{-\frac{1}{2}} \newline
L(y) = \int_0^{12}\sqrt{1+\frac{1}{2x}}dx=\frac{1}{2}(ln{5+2\sqrt{6}}+10\sqrt{6})=13.3937"

"\int_0^{12}\sqrt{1+\frac{1}{2x}}dx"

apply substitution "u=2x"

"=\int_0^{24}\frac{\sqrt{u+1}}{2\sqrt{u}}du"

take the constant out "\int{a f(x)}dx=a\int{f(x)}dx"

"=\frac{1}{2}\int_0^{24}\frac{\sqrt{u+1}}{\sqrt{u}}du"

apply integration by parts "u=\frac{\sqrt{u+1}}{\sqrt{u}}, v'=1"

"=\frac{1}{2}[\sqrt{\frac{u+1}{u}}u-\int-\frac{1}{2u^{\frac{1}{2}}\sqrt{u+1}}du]_0^{24} \newline
\int-\frac{1}{2u^{\frac{1}{2}}\sqrt{u+1}}du=-ln|\sqrt{u+1}+\sqrt{u}| \newline
=\frac{1}{2}[\sqrt{\frac{u+1}{u}}u-(-ln|\sqrt{u+1}+\sqrt{u}| )]_0^{24}"

Simplify

"=\frac{1}{2}[\sqrt{\frac{u+1}{u}}u+ln|\sqrt{u+1}+\sqrt{u}|]_0^{24}"

Compute the boundaries:

"ln(5+2\sqrt{6}) + 10\sqrt{6}\newline
=\frac{1}{2}(ln(5+2\sqrt{6}) + 10\sqrt{6})=13.3937"